T(n) = 2T(n/2) + n log log n

T(n) = 2T(n/2) + n log log n
Answer #1

Master Theorm - T(n) = a T (h/b) + 0 (nh log in) a=1, b>, KJO and f is real number. Dif as bk, Then Ton) = 0 (alog ) 2) if a=now, given that T (n) = 2T(1/2) + n log logn Tin) = 2T (/) to cnlogen) a=2, 6-12, K=O, D=2 b = 2 = 2 : a = b k and Ps-1 u. T